rebeccaxhawaii (rebeccaxhawaii):

*boring story that tries to make math interesting* To open these doors, you must speak three functions in standard form. One function, f(x), with two real rational solutions. One function, g(x), with two real irrational solutions. One function, h(x), with two complex solutions.

rebeccaxhawaii (rebeccaxhawaii):

@johnweldon1993

rebeccaxhawaii (rebeccaxhawaii):

okay so this is only the first question

OpenStudy (johnweldon1993):

....I wanna know the story >.< XD

rebeccaxhawaii (rebeccaxhawaii):

if you must ... For this adventure, you and world renowned Professor Sherlock McMerlock are traveling to the Lost Island of Laplaya. When the boat arrives on the island shore, you and Professor McMerlock disembark on your adventure. You trudge through the jungles and arrive at three impressively large doors. About eye level on each door is an intricately carved keyhole. Directions are scratched into the wood above each keyhole.

OpenStudy (johnweldon1993):

...It's SO GOOD!!! Lol okay..so It relates to the quadratic formula, more specifically the discriminant \[\large \sqrt{b^2 - 4ac}\] If \(\large b^2 - 4ac\) = a perfect square...we have 2 rational number solutions if \(\large b^2 - 4ac\) does not = a perfect square, we have 2 irrational number solutions and if \(\large b^2 - 4ac\) is less than 0...we get 2 complex solutions

OpenStudy (johnweldon1993):

So we just need to come up with some function \(\large ax^2 + bx + c\) where \(\large b^2 - 4ac\) fits those criteria

OpenStudy (owlcoffee):

You can define any type of function with a standard form, let it be polynomial or exponential. For example, the first function can be represented as a polynomial in order to keep it simple, remember that rational are part of the real numbers, rational. So we can plot any rational number on the structure: \[f(x)=(x-a)(x-a_2)...(x-a_n)\] where all those "a"s represent the roots or the "solutions" to the function. I'll take, for instance two numbers: \(\frac{ 5 }{ 4 }\) and \(-\frac{ 10 }{ 7 }\) which adds some variation: \[f(x)=(x-\frac{ 5 }{ 4 })(x-(-\frac{ 10 }{ 7 })) \iff f(x)=(x-\frac{ 5 }{ 4 })(x + \frac{ 10 }{ 7 })\] You can apply the distributive if you want to find the general form of that function.

rebeccaxhawaii (rebeccaxhawaii):

so i found some answers would f(x)= x^2+6x+16 , g(x)= x^2+6x+1 , h(x)+2x^2+4x+5 would this work ?

OpenStudy (johnweldon1993):

f(x) \(\large 6^2 - 4(1)(16) = 36 - 72 = -36\) Nope...to be real...it has to be > 0

rebeccaxhawaii (rebeccaxhawaii):

wtf

OpenStudy (johnweldon1993):

f(x) and g(x) both have to be real...so those discriminantes have to be > 0

rebeccaxhawaii (rebeccaxhawaii):

i thought g(x) doesnt have to be

OpenStudy (owlcoffee):

But what assures the solutions will be rational? They will for sure be real, but consider that the solution might as well be irrational.

OpenStudy (johnweldon1993):

One function, f(x), with two ****real**** rational solutions. > 0 One function, g(x), with two ****real**** irrational solutions. > 0 One function, h(x), with two ****complex**** solutions. < 0

OpenStudy (owlcoffee):

Yes, with that they mean a real number structured as a rational and irrational, else she would just put a "2" and it would solve the problem.

rebeccaxhawaii (rebeccaxhawaii):

ok i gotcha but what do you mean by less than 0

OpenStudy (johnweldon1993):

The quadratic formula is \[\large \frac{-b \pm \sqrt{b^2 - 4ac}}{2a}\] Remember how I said the discriminant is that \(\large b^2 - 4ac\) part? Well see how it is under the square root sign? if the discriminant comes out negative...we take the square root of a negative number....hence it would be complex solutions

rebeccaxhawaii (rebeccaxhawaii):

john john john .. can you talk rebecca style please

OpenStudy (johnweldon1993):

XDDDD okay..lets see

OpenStudy (johnweldon1993):

Let me give you an example, maybe that will help

rebeccaxhawaii (rebeccaxhawaii):

okay okay

OpenStudy (johnweldon1993):

So g(x) has to be 2 real irrational solutions \[\large ax^2 + bx + c\] I need some a,b and c that make \(\large b^2 - 4ac\) a number > 0 so hmm...lets do b = 6 , a = 1 and c = 2 *just completely random numbers* let see if it works? \[\large b^2 - 4ac \rightarrow 6^2 - 4(1)(2) \rightarrow 36 - 8 \rightarrow 28\] Since that is a number > 0 and we know 28 is NOT a perfect square \(\large x^2 + 6x + 2\) Will have 2 real irrational numbers

OpenStudy (johnweldon1993):

Tell me ANYTHING that doesn't make sense

rebeccaxhawaii (rebeccaxhawaii):

this might take a while to process this as you know im quite slow

OpenStudy (johnweldon1993):

Lol I got time :)

OpenStudy (johnweldon1993):

Actually...it might help reading about it elsewhere as well? http://www.regentsprep.org/regents/math/algtrig/ate3/discriminant.htm

rebeccaxhawaii (rebeccaxhawaii):

john im done with school in just a bit do you think we can continue tomorrow

OpenStudy (johnweldon1993):

Okay, sorry if I hurt your head :/ </3 XD

rebeccaxhawaii (rebeccaxhawaii):

you didnt its the math thats killing me ... slowly ... like my brain

rebeccaxhawaii (rebeccaxhawaii):

any who get good rest youll need it

OpenStudy (johnweldon1993):

Lol take a breather...get a snack...and some juice...but share :)

rebeccaxhawaii (rebeccaxhawaii):

@johnweldon1993

OpenStudy (johnweldon1993):

Oh back to this story!! Almost forgot it, gave me chills XD Okay first, I want you to read that link I posted above :)

rebeccaxhawaii (rebeccaxhawaii):

okay sorry im reading it

OpenStudy (johnweldon1993):

Lol it's okay...math cheating on me it's cool ;P

rebeccaxhawaii (rebeccaxhawaii):

actually the class is having time of prayer so yeah i guess they are similar lol

OpenStudy (johnweldon1993):

XDDDDD lol *foot in mouth* XDDD

rebeccaxhawaii (rebeccaxhawaii):

okay i think i got it

OpenStudy (johnweldon1993):

Okay...so I want a quadratic with 2 real IRRATIONAL solutions...

rebeccaxhawaii (rebeccaxhawaii):

what is that

OpenStudy (johnweldon1993):

Lol your original question "One function, f(x), with two real rational solutions. One function, g(x), with two real irrational solutions. One function, h(x), with two complex solutions." So we need those...based on what you read...how can we do the g(x) one?

rebeccaxhawaii (rebeccaxhawaii):

when we plug it in and check it it will work but idk how to get a problem like that

OpenStudy (johnweldon1993):

Okay...so based on what you read...if we have some function \(\large ax^2 + bx + c\) where the discriminant is \(\large b^2 - 4ac\) If the discriminant is greater than 0 and not a perfect square...we will have our g(x) satisfied So we just pick random numbers for a,b and c until we make that work

rebeccaxhawaii (rebeccaxhawaii):

REBECCA WORDS PLS

OpenStudy (johnweldon1993):

Lol no these are JOHN WORDS!!! But i mean, you should get that though, based on the link you read

rebeccaxhawaii (rebeccaxhawaii):

expecting me to read johnny dont expect miracles

rebeccaxhawaii (rebeccaxhawaii):

jk ima look again

OpenStudy (johnweldon1993):

Remember my example I did before I chose random numbers for a,b and c where after calculating \(\large b^2 - 4ac\) I got a number greater than 0 that wasnt a perfect square...so it would have worked for g(x) I'll repost it So g(x) has to be 2 real irrational solutions \[\large ax^2+bx+c\] I need some a,b and c that make \(\large b^2−4ac\) a number > 0 so hmm...lets do b = 6 , a = 1 and c = 2 *just completely random numbers* let see if it works? \[\large b^2−4ac→6^2−4(1)(2)→36−8→28\] Since that is a number > 0 and we know 28 is NOT a perfect square \(\large x^2+6x+2\) Will have 2 real irrational numbers

rebeccaxhawaii (rebeccaxhawaii):

wo complex roots ?

rebeccaxhawaii (rebeccaxhawaii):

two *

OpenStudy (johnweldon1993):

Are you saying you want to do that next? Because that is not what I just did above...that was for the 2 REAL irrational roots

rebeccaxhawaii (rebeccaxhawaii):

IM SO CONFUSED

OpenStudy (johnweldon1993):

Okay forget everything so far Focus on this equation \(\large b^2 - 4ac\) Give me 3 random numbers that I can plug in for a,b and c

rebeccaxhawaii (rebeccaxhawaii):

this is my type of learning

rebeccaxhawaii (rebeccaxhawaii):

okay 593

OpenStudy (johnweldon1993):

Lol I'm gonna get you through this hun :P Okay so you're giving me 5 for a 9 for b and 3 for c right?

rebeccaxhawaii (rebeccaxhawaii):

indeedy

OpenStudy (johnweldon1993):

Okay So with a = 5, b = 9 and c = 3 From this equation \(\large b^2 - 4ac\) that would give me \[\large 9^2 - 4(5)(3) = 81 - 60 = 21\] Good so far?

rebeccaxhawaii (rebeccaxhawaii):

yes

OpenStudy (johnweldon1993):

Okay...now comes the learning part When \(\large b^2 - 4ac\) turns out to be greater than 0...and not a perfect square, then we would get 2 real irrational roots Dont worry about why for now....just tell me if you accept that :)

rebeccaxhawaii (rebeccaxhawaii):

wait so how is it a perfect square ?

rebeccaxhawaii (rebeccaxhawaii):

i understand everything else

OpenStudy (johnweldon1993):

Tell me what a perfect square is :)

rebeccaxhawaii (rebeccaxhawaii):



OpenStudy (johnweldon1993):

XD you're perfect XD

OpenStudy (johnweldon1993):

Okay...but should i take that to mean you're not sure what a perfect square number is?

rebeccaxhawaii (rebeccaxhawaii):

64

OpenStudy (johnweldon1993):

Okay good...and the reason is because we know 8^2 = 64 right?

rebeccaxhawaii (rebeccaxhawaii):

yes

OpenStudy (johnweldon1993):

Perfect, we're getting there :) So in the numbers you gave me...and I went through the equation and got 21 as an answer...is 21 a perfect square?

rebeccaxhawaii (rebeccaxhawaii):

no its not

OpenStudy (johnweldon1993):

Right!!! So, we went through the equation...and got a number GREATER than 0....AND it was NOT a perfect square....from that link I gave you....that means we will have 2 real irrational roots

rebeccaxhawaii (rebeccaxhawaii):

okay yay i get it so far

OpenStudy (johnweldon1993):

Okay good...lets finish that one up You gave me a = 5, b = 9 and c = 3 Now we just need to plug those into the equation for the actual function...because that \(\large b^2 - 4ac\) was just for the first part...now the second part is using the equation \(\large ax^2 + bx + c\) we just need to pug in the a,b and c you gave me So the function would be \(\large 5x^2 + 9x + 3\)

rebeccaxhawaii (rebeccaxhawaii):

yes

OpenStudy (johnweldon1993):

Okay good! that's 1 done......we have g(x) done Now lets work on f(x) So again...give me 3 random numbers

rebeccaxhawaii (rebeccaxhawaii):

im gonna write this down

OpenStudy (johnweldon1993):

Okay, I'll hang on for a bit :)

rebeccaxhawaii (rebeccaxhawaii):

so what after a bit you just gonna leave okay i see how it is

OpenStudy (johnweldon1993):

Lol haha I was giving you time to write it down you butt!! ;P

rebeccaxhawaii (rebeccaxhawaii):

hah jk but my numbers are 274

OpenStudy (johnweldon1993):

;P Okay so lets try it out a=2 b=7 c=4 So now...after we go through the \(\large b^2 - 4ac\) we WANT a number greater than 0...but we also want a perfect square this time...because that will give us 2 real rational solutions like we want for fx) So lets try it \(\large 7^2 - 4(2)(4) = 49 - 32 = 17\) sadly not a perfect square...so we need new numbers :)

rebeccaxhawaii (rebeccaxhawaii):

924

OpenStudy (johnweldon1993):

Okay a=9 b=2 c=4 \[\large 2^2 - 4(9)(4) = 4 - 144 = -140\] Oh oh...okay wait forget f(x) for a sec...see how we came up with a negative number??

rebeccaxhawaii (rebeccaxhawaii):

no i mean 294

OpenStudy (johnweldon1993):

Lol okay fine...we'll come back to the above though.... Okay so a = 2 b = 9 c = 4 \[\large 9^2 - 4(2)(4) = 81 - 32 = 49\] and is 49 a perfect square??

rebeccaxhawaii (rebeccaxhawaii):

hold you applause everyone i did that to make it like that thank you thank you

OpenStudy (johnweldon1993):

Hahahahahaha I almost missed that you replied because i was still clapping ;)

rebeccaxhawaii (rebeccaxhawaii):

sweet as a lemon

rebeccaxhawaii (rebeccaxhawaii):

so i win right ?

OpenStudy (johnweldon1993):

Lol you win for f(x) yes...we found 3 numbers that make the discriminant a perfect square greater than 0 So \(\large f(x) = 2x^2 + 9x + 4\)

OpenStudy (johnweldon1993):

Now we just need h(x) And you have already done it actually...remember where I said "we're coming back to this one"